OpenStudy (readergirl12):

3 consecutive natural numbers such as the sum of the squares is 7502. What is the first of these three numbers ? I'm not sure if this is 8th grade math...I don't even understand the question.. ;-;

Directrix (directrix):

Natural numbers are the counting numbers {1, 2, 3, 4, ........} Consecutive natural numbers (think of them as positive integers) look like this: 2, 3, 4 or maybe 60, 61, 62. Note that two consecutive integers differ by one.

Directrix (directrix):

In your problem, let the consecutive natural numbers be: x, x+1, and x + 2.

OpenStudy (welshfella):

let the natural numbers be x , x+1 and x+ 2

Directrix (directrix):

When you square them, they look like this: x^2 ( x+1)^2 , and (x + 2^2.

Directrix (directrix):

Sum of the squares means to add them. x^2 + ( x+1)^2 + (x + 2) ^2 = 7502

Directrix (directrix):

The task is to solve this equation for x: x^2 + ( x+1)^2 + (x + 2) ^2 = 7502

Directrix (directrix):

Multiply out the squared expressions and then simplify. Post what you get, okay?

Directrix (directrix):

( x+1)^2 = ? (x + 2) ^2 = ? @readergirl12

Directrix (directrix):

I will help you do this if you will help out on the multiplication. We can do the problem in parts and then put them together to get the answer.

OpenStudy (readergirl12):

Um..Does the "^" mean squaring?

Directrix (directrix):

Yes.

Directrix (directrix):